In an article about investment alternatives, Money Magazine reported that drug stocks provide a potential for long-term growth, with over fifty percent of the adult-population of the USA taking prescription drugs on a regular basis. For adults 65 and older, 82% take prescription drugs regularly. For adults age 18 to 64, 49% take prescription drugs regularly. The 18-64 group accounts for 83.5% of the adult population (Statistical Abstract of the United Staes, 2008).

a) GIVEN that an adult takes prescription drugs regularly, what is the probability that the adult is 65 or older?
Answer #1

Given that The drug stocks provi de a potential for long-term growth, with over 50% of the adult population of the United Sta
Answer #2
a. Adults are made up of those 18-64 and those 65+
Thus,P(an adult is 65+) = 1 - P(an adult is 18-64) =
1 - .835 = .165 or 16.5%

b.For this problem, we need to figure the P(adult is 65 or older | the adult takes drugs) = P(adult 65 or older and takes drugs)/P(adult takes drugs)

The numerator equals .165 * .82
The denominator equals .165 * .82 + .835 * .49 (including those 65+ and those 18-64 taking drugs)

Then, P(adult is 65 or older | the adult takes drugs) =
.165 * .82/(.165 * .82 + .835 * .49 ) = 0.248507668 = .249 = 24.9%.
Answer #3
a. Adults are made up of those 18-64 and those 65+
Thus,P(an adult is 65+) = 1 - P(an adult is 18-64) =
1 - .835 = .165 or 16.5%

b.For this problem, we need to figure the P(adult is 65 or older | the adult takes drugs) = P(adult 65 or older and takes drugs)/P(adult takes drugs)

The numerator equals .165 * .82
The denominator equals .165 * .82 + .835 * .49 (including those 65+ and those 18-64 taking drugs)

Then, P(adult is 65 or older | the adult takes drugs) =
.165 * .82/(.165 * .82 + .835 * .49 ) = 0.248507668 = .249 = 24.9%.